On September 1, 2013, Susan Chao bought a motorcycle for $36,000. She paid $1,000 down and financed the balance with a five-year loan at an APR of 7.8 percent, compounded monthly. She started the monthly payments exactly one month after the purchase (i.e., October 1, 2013). Two years later, at the end of October 2015, Susan got a new job and decided to pay off the loan.

  

If the bank charges her a 1 percent prepayment penalty based on the loan balance, how much must she pay the bank on November 1, 2015?
Answer #1

Cost of motorcycle = $36,000

Down payment = $1,000

Loan amount = $36,000 - $1,000 = $35,000

Term of loan = 5 years

Number of monthly loan payments = n = 5 years * 12 = 60

APR = 7.8%

Monthly interest rate = r = 7.8%/12 = 0.65% = 0.0065

Loan payment = Loan amount*r / {1-(1+r)-n}

Monthly loan payment = ($35,000*0.0065)/(1-1.0065-60) = $706.33

Loan balance outstanding = Present value of remaining loan payments

Present value of annuity = Annuity amount*{1-(1+r)-n}/r

Number of payments paid in 2 years = 2 years * 12 = 24

Number of loan payments remaining after 2 years = 60 – 24 = 36

Loan balance outstanding as on November 1, 2015 = $706.33*(1-1.0065-36)/0.0065 = $22,606.72

Prepayment penalty = $22,606.72 * 1% = $226.07

Total amount paid to bank = $22,606.72 + $226.07 = $22,832.79